Fermi National Laboratory
 

Operation Summary for February 11, 2011 - February 14, 2011

 

The Integrated Luminosity for the period from 2/7/11 to 2/14/11 was 52.78 Pb-1 (inverse picobarns). NuMI reported receiving 7.33E18 protons on target during this same period.

Friday Feb. 11

The day shift began with Operations monitoring the Tevatron (TEV) store 8487 (50.08E30) and the Recycler stash (291.7E10), with the Antiproton Source (Pbar) stacking (25.1E10), with the Linac, Booster, and Main Injector (MI), operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8487 at 12:50 PM, put the TEV into HEP shot setup, and established store 8488 at 2:08 PM with an average initial luminosity (AIL) of 350.3E30. Operations disabled the MiniBooNE beam at 10:53 AM due to a frozen sump pump discharge line that required an access by FESS personnel to fix. Shortly after 3 PM, a beam movement in the TEV resulted in a slight loss of luminosity; operators investigated and found the cause was an earthquake in Chile.

During the evening shift, FESS finished the MiniBooNE sump repairs and Operations had beam restored by 7:30 PM. At 8:40 PM, Operations reported the loss of store 8488 due to an abort aused by the TEV ramp controller power supply. Operators investigated, but found no reason for the abort. Operators prepared the D0 collision hall for an access. Operations reported that a MI RF station (MIRF4) had tripped off multiple times. Operators began the TEV turn on at the end of the shift.

Saturday Feb. 12

The midnight shift began with Operations monitoring the stash (278.8E10), with Pbar stacking (13.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators recovered the TEV and put it into shot setup at 2:56 AM. Operations established store 8492 at 4:10 AM with an AIL of 322.55E30. Operations reported that four Linac low energy RF stations (1-4) tripped off multiple times during the shift.

During the day shift, at 1:40 PM, an off-site ComEd power glitch caused the loss of the store, the Recycler antiproton stash, many cryo systems, and the Pbar Accumulator antiproton stack. The glitch caused failures in equipment throughout the accelerator chain. The remainder of the shift was spent recovering the accelerators. Numerous personnel came in to assist. There was no beam delivered for the rest of the shift.

During the evening shift, operators started the shift still recovering the accelerators. Operations restored beam to NuMI at 9 PM and to MiniBooNE at 9:36 PM. Pbar personnel accessed the Pbar enclosure to repair a shunt power supply. Pbar resumed stacking at 11:37 PM.

Sunday Feb. 13

The midnight shift began with Operations monitoring the Accelerator Complex, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that Cryo System personnel had the TEV cryo system restored by midnight. The TEV turn on and check out lasted most of the rest of the shift. Operators had the TEV in stand-by mode at the end of shift. CDF personnel found and fixed several problems with their detector: CDF controls issues, power supply, and a power supply ramp card. The glitch hit the Pelletron hard. Several Pelletron personnel came in and accessed to recover the Pelletron. Multiple power supplies needed to be replaced and problems persisted until the end of the shift. Operations reported several more MI and Linac RF trips.

During the day shift, Pelletron personnel reported at 9:11 AM that they had recovered the Pelletron. Operations put the TEV into shot setup and established store 8497 at 10:33 AM with an AIL of 162.49E30. Operations tuned on the Proton Source and reported minor down times caused by Linac and Booster RF stations.

During the evening shift, Operations monitored the Accelerator Complex.

Monday Feb. 14

The midnight shift began with Operations monitoring store 8497 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8497 at 1:33 AM, put the TeV into shot setup, and established store 8498 at 2:56 AM with an AIL of 328.81E30. At 6:40 AM, the MI safety system went down due to a ground fault indication and caused the loss of ~135E10 antiprotons in the Recycler. Interlock technicians began investigating at the end of shift. Operations reported minor down times caused by Linac and Booster RF stations.

The Plans for Monday

The plans for today are to bring MI back on line and resume stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/14/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory